The area between z = 0 and z = 2 is

1
Expert's answer
2022-03-10T09:00:48-0500

Area between "z=0" and "Z=2"

"P(0<Z<2)=P(Z=2)-P(Z=0)"

From z distribution table

"=0.9772-0.500=0.4772"